# Large Language Models as Realistic Microservice Trace Generators

### v1

@@ Abstract
Computer system workload traces, which record hardware or software events during application execution, are essential for understanding the behavior of complex systems and managing their processing and memory resources. However, obtaining real-world traces can be challenging due to the significant collection overheads in performance and privacy concerns that arise in proprietary systems. As a result, synthetic trace generation is considered a promising alternative to using traces collected in real-world production deployments. This paper proposes to train a large language model (LLM) to generate synthetic workload traces, specifically microservice call graphs. To capture complex and arbitrary hierarchical structures and implicit constraints in such traces, we fine-tune LLMs to generate each layer recursively, making call graph generation a sequence of easier steps. To further enforce learning constraints in traces and generate uncommon situations, we apply additional instruction tuning steps to align our model with the desired trace features. Our evaluation results show that our model can generate diverse realistic traces under various conditions and outperform existing methods in accuracy and validity. We show that our synthetically generated traces can effectively substitute real-world data in optimizing or tuning systems management tasks. We also show that our model can be adapted to perform key downstream trace-related tasks, specifically, predicting key trace features and infilling missing data given partial traces. Codes are available in https://github.com/ldos-project/TraceLLM.
@@ 1 Introduction
Computer system workload traces document hardware or software events that occur as applications execute on computing machines, receive requests, process them, and serve responses. Such traces are vital for analyzing complex computer systems and optimizing their CPU, memory, IO and networking resource allocation and management. However, obtaining real-world traces is often hindered by privacy concerns and their general unavailability. As an alternative, synthetic traces provide limitless size and variety, offering significant advantages for testing and analysis, including the ability to simulate challenging conditions like stress-testing environments. While recent advances in generative machine learning, including LSTMs (Sherstinsky, 2020), GANs (Goodfellow et al., 2014), and diffusion models (Ho et al., 2020), have improved synthetic trace generation, these methods typically only generate specific fields, such as the number of requests or resource types (Bergsma et al., 2021), or are confined to fixed-structure traces, like network packets (Jiang et al., 2023; Yin et al., 2022).
@@
In this paper, we show how to use large language models (LLMs), transformer-based (Vaswani et al., 2017) neural networks pre-trained autoregressively on large and diverse text datasets (Brown et al., 2020; Touvron et al., 2023), to generate synthetic workload traces. It has been shown that LLMs can be readily adapted to model a variety of domains besides natural language, such as protein sequences (Shen et al., 2024), code (Roziere et al., 2023), and tabular data (Borisov et al., 2023). LLMs can produce outputs that are well-aligned with user inputs in several flexible ways such as fine-tuning model weights (Ouyang et al., 2022; Wei et al., 2021), and can generalize to new user inputs at inference (Chung et al., 2022; Sanh et al., 2021). Thus, LLMs have the potential to generate synthetic traces that accurately model the structure of real-world traces while following user prompts.
@@
Despite their potential, using LLMs for synthetic trace generation presents significant challenges. Traces are often logged in a tabular format and follow an underlying structure such as a graph, where the graph can be arbitrarily deep and wide. This means that it is non-trivial to represent valid traces as text sequences, which is the format best suitable for modern autoregressive LLMs. Moreover, there are often complex implicit constraints in trace data that rely on relationships between multiple trace features. For example, an application process’s start time must be earlier than the start time of all the child processes it spawns, while the parent process’s end time must be later than the end time of its children; such constraints need to hold at all nodes in the application’s graphical representation.
@@
Our proposed approach adapts general-purpose LLMs to generate synthetic system traces, with a particular focus on microservice call graphs, a type of trace with rich directed acyclic graph (DAG) structures. We represent these graphs in a text-based format suitable for LLMs, enabling their use in trace generation. One of our key innovations is to generate call graphs with structural constraints by recursively generating subgraphs, or layers. This approach allows the model to break down the complex task of reasoning about hierarchical graph structures and complex constraints into multiple easier tasks. To further enhance the model’s ability to follow structural constraints and to meet user-requested attributes, we perform instruction tuning. During this phase, the model learns to explicitly generate a series of intermediate instructions between recursive layer generation steps, performing simple arithmetic and logical checks to ensure adherence to the desired structure.
@@
We demonstrate the effectiveness of our approach by fine-tuning Llama-2 7B (Touvron et al., 2023) with our method on microservice trace data and performing a series of evaluations. The results show that the proposed recursive generation and use of intermediate instructions significantly enhance the model’s ability to generate valid outputs for complex (i.e., deep and wide) call graph structures. When compared to traces from learned generative models and a probabilistic model, synthetic traces produced by our model more closely match the distribution of real traces. We further show that our fine-tuned model performs well when used for downstream tasks for trace feature prediction compared to Llama-3.1 405B, one of the state-of-the-art LLMs.
We summarize our key contributions below:
  • $\bullet$
    We introduce a novel method for creating valid and rich synthetic microservice traces using LLMs: (1) We recursively generate layers of subgraphs along with prompts for subsequent layers and (2) We also align the LLM to describe the generated layers with intermediate instructions.
  • $\bullet$
    We show that recursive generation and intermediate instructions improve the validity of synthetic traces. Also, synthetic traces generated by our model are more realistic regarding distribution similarities than those from a baseline generative model and a handcrafted expert model.
  • $\bullet$
    We demonstrate that synthetic traces can effectively substitute real data for training microservice management tasks. Furthermore, our model generalizes to unseen user-requested attributes during inference and can be further adapted for downstream tasks like infilling missing trace data.
@@ 2 Background
Figure 1: A simple social network application consists of eight microservices (Huye et al., 2023). Each user request triggers a sequence of microservice calls, forming a microservice call graph. The red lines represent the microservice call graph for a user request. Microservice call graphs are commonly logged in a tabular format, as shown in the figure on the right. Each row in the table represents a communication between two microservices, with the details of the communication logged as features in the columns.
@@
Microservice Call Graphs.  In modern software architecture, an application is typically constructed as a constellation of multiple microservices (Gan et al., 2019; Luo et al., 2022; Huye et al., 2023), each with specific functionalities and dependencies on one another. When users interact with these applications, for instance, by sending HTTP requests to web servers, a complex sequence of communications among these microservices is triggered. Thus, a user request induces a microservice call graph, which maps the control/data flow and dependencies among the microservices involved in fulfilling the user’s request.
@@
Figure 1 is an example of a social network application deployed with several microservices (8 in total). In the figure, the red arrows indicate communications between microservices involved in processing the user’s request. The request is sent to a microservice (e.g., “Front end” in Figure 1) and waits for the communication to terminate. If the microservice requires additional communication to handle the request, then it triggers another microservice call (e.g., from “Front end” to “Authentication” in Figure 1). These communications triggered by a user’s request form a microservice call graph with four microservices. The vertices of the graph correspond to microservices (or the client), while the edges correspond to API calls invoking the microservices. Note that some edges are not part of the call graph as the corresponding microservices are not invoked in processing this particular request.
@@
Each call graph can be represented as a tabular log trace with a textual description of the features of each API call (i.e., edges), including the source and destination of the request, type of request (e.g., HTTP and RPC), and start/finish time. As call graphs have a hierarchical structure, the tabular trace should preserve the parent-child relationships by ensuring that the child’s source matches the parent’s destination. Moreover, the start and end times of each call should be consistent with each other: (1) the start time of a microservice must be earlier than its finish time, and (2) the parent-child relationships must be honored, i.e., the parent’s start time must precede the child’s, and parent’s finish time must follow the child’s. Finally, the IDs within a call graph (dot-decimal numbers provided for each call) must also be hierarchically connected to form a DAG structure.
@@
Synthetic Trace Generation using Machine Learning.  The analysis of microservice traces plays a pivotal role in improving the performance and reliability of services, and guides techniques that enable high-performance and efficient use of the underlying machines. Representative use cases include critical path analysis (Zhang et al., 2022), anomaly detection (Xie et al., 2023), root cause analysis (Ikram et al., 2022), cluster management (Qiu et al., 2020), and cluster scheduling (Singhvi et al., 2021). Unfortunately, access to traces remains challenging due to business and privacy concerns.
@@
Given the importance and limited availability of public computer system traces, including microservice traces, several recent studies have explored generative models for synthetic trace generation. Existing works (Lin et al., 2020; Jiang et al., 2023) leverage GAN (Goodfellow et al., 2014) and diffusion (Ho et al., 2020) models to generate network packet traces, while other work (Bergsma et al., 2021) uses LSTMs (Sherstinsky, 2020) to generate virtual machine workload traces. Even though the generative models have shown effectiveness in each domain, the methods are used only for predicting specific fields or following training data distributions without conforming to structural constraints. These methods do not apply to microservice call graphs because they cannot handle the hierarchical structures of the call graphs.
@@
Since traces have specific structures that can be represented in tabular form, machine learning methods for synthetic tabular data generation could be applied to synthetic trace generation. Recent approaches, such as TVAE (Xu et al., 2019) and GReaT (Borisov et al., 2023), leverage VAE (Kingma & Welling, 2013) and language models to advance synthetic tabular data generation techniques. However, these methods have limitations when applied to microservice traces, as they do not account for the hierarchical structure of call graphs within tabular representations. We provide a detailed comparison with tabular data generation methods in § 4.
3 Training LLMs to Generate Microservice Traces
@@
Our goal is to train a generative model for microservice call graph traces. We want to allow end-users to simulate various scenarios, such as stress-testing a novel software application or feature, by conditioning the model’s output on user-requested attributes including the application being invoked, the number of microservice communications (i.e., graph edges), and the overall latency of the application. Given the limitations of existing trace generation approaches, we turn to LLMs, which are transformer-based (Vaswani et al., 2017) models with billions of parameters. We initialize our model from a general-purpose LLM pre-trained on a large and diverse text dataset, as these models have shown effectiveness when adapted for specialized domains such as proteins (Shen et al., 2024) and code (Roziere et al., 2023). In addition, LLMs can be conditioned in a variety of arbitrary manners, including natural language prompting (Ouyang et al., 2022) and structured input sequences (Borisov et al., 2023).
This section presents our approach for training an LLM to generate microservice call graphs. We train our model in two stages. In the first stage, we pre-train the model to learn the complex interactions between the vertices and edges in real call graph data. We describe how we encode call graphs, stored as tabular data, into a text format that can be tokenized and processed by the LLM. Then, we detail a novel approach to improve the model’s generation of large, complex graph structures. We propose to decompose the graph generation task into a series of simpler, recursive subgraph generation tasks that allow the model to reason about local features while respecting global structure. In the second training stage, we fine-tune the model to follow user instruction and allow flexible generation of call graphs with desired attributes. During instruction tuning, we propose to include a series of natural language reasoning steps that reinforce the model’s ability to adhere to constraints during inference.
@@ 3.1.1 Encoding Call Graphs as Text
LLMs expect sequences of text as input, so we must encode our dataset of call graphs into text-based representations before training our model. As detailed in § 2 and shown in Figure 1, microservice call graphs are initially stored as tables. Rows represent edges (i.e., communications between microservices), while columns describe features for each edge. We follow the method proposed by GReaT (Borisov et al., 2023) and encode features in a natural language format, which requires minimal pre-processing and allows us to exploit the LLM’s pre-training on diverse, natural texts. Our lossless encoding procedure preserves all necessary information to recover the unique graph that produced the tabular data. Besides edge features, we also encode global attributes of the call graph to serve as conditioning information for the model.
Tabular call graph $\mathbf{X}$ has $m$ columns of features with textual names $\{f_{1},f_{2},\dots,f_{m}\}$ and $n$ rows of edges $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}$. We denote the value of feature $j\in\{1,\dots,m\}$ for edge $i\in\{1,\dots,n\}$ as $v_{ij}$. We encode each edge $\mathbf{x}_{i}$ as a text sequence $\mathbf{t}_{i}=[t_{i1},t_{i2},\dots,t_{im}]$, where $t_{ij}$ is a description of the $j$-th feature with the format $t_{ij}=[\phi(f_{j}),v_{ij}]$. Here, $\phi(f)$ encodes feature name $f$ into a text template with a subject-predicate structure to provide a natural language description of feature value $v_{ij}$. For instance, the encoding for edge 1 in Figure 1 would be [Edge ID is 0, Source is Client, Destination is Front end, Type is HTTP, Communication starts at 0 ms, Communication finishes at 24 ms]. We encode tabular call graph $\mathbf{X}$ to the equivalent text-based representation $\mathbf{t}=[\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{n}]$, formed as a sequence of the text-encoded edges $\mathbf{t}_{i}$. We note that the structure and constraints of the call graph only depend on the feature values and are invariant to the specific feature ordering imposed by the columns of the tabular data. Therefore, during training, we randomly shuffle the order of the features within each edge as in Borisov et al. (2023) to remove any spurious associations that arise from position information.
Apart from individual edges, the overall call graph can also be described by attributes, including the maximum depth, the total number of edges, and the total communication latency. These attributes are useful for summarizing the complex interactions between edges and can be fed to the model as a prompt to condition call graph generation. Let call graph $\mathbf{X}$ have $r$ attributes with names $\{a_{1},a_{2},\dots,a_{r}\}$ and corresponding values $\{w_{1},w_{2},\dots,w_{r}\}$. We encode the attributes as a text sequence $\mathbf{c}=[c_{1},c_{2},\dots,c_{r}]$, where $c_{j}$ is a description of the $j$-th attribute with the format $c_{j}=[a_{j},``:",w_{j}]$. See the Conditions shown in red in Figure 2 for a simplified example of text-encoded call graph attributes. We include the attributes at the start of each text-encoded call graph sequence and predict them along with the edge tokens during pre-training. Similar to the edge features, we randomly shuffle the order of graph attributes during training. We additionally drop each attribute independently with probability $p_{drop}$ to allow flexible prompting with arbitrary subsets of attributes.

### v2

@@ Abstract
Workload traces are essential to understand complex computer systems’ behavior and manage processing and memory resources. Since real-world traces are hard to obtain, synthetic trace generation is a promising alternative. This paper proposes a first-of-a-kind approach that relies on training a large language model (LLM) to generate synthetic workload traces, specifically microservice call graphs. To capture complex and arbitrary hierarchical structures and implicit constraints in such traces, we show how to fine-tune LLMs to generate recursively, making call graph generation a sequence of easier steps. To further enforce learning constraints in traces and generate uncommon situations, we argue for applying additional instruction tuning steps to align our model with the desired trace features. Our evaluation results show that we can generate diverse realistic traces under various conditions and outperform existing methods in accuracy and validity. We demonstrate that our synthetically generated traces can effectively replace real data to optimize important microservice management tasks. Additionally, our model adapts to downstream trace-related tasks, such as predicting key trace features and infilling missing data.
Large Language Models as Realistic Microservice Trace Generators
Donghyun Kim, Sriram Ravula, Taemin Ha, Alexandros G. Dimakis, Daehyeok Kim, Aditya Akella University of Texas at Austin {donghyun,sriram.ravula,taemin.ha,dimakis}@utexas.edu, {daehyeok,akella}@cs.utexas.edu
@@ 1 Introduction
Computer system workload traces document hardware or software events that occur as applications execute on computing machines, receive requests, process them, and serve responses. Such traces are vital for analyzing complex computer systems and optimizing their CPU, memory, networking resource allocation, and management. However, obtaining access to real-world traces is often challenging due to limited public data availability and the difficulty of collecting them at large scale from diverse environments, especially in complex cloud computing settings. As an alternative, synthetic traces provide limitless size and variety, offering significant advantages for testing and analysis, including the ability to simulate challenging conditions like stress-testing environments. While recent advances in generative machine learning, including LSTMs (Sherstinsky, 2020), GANs (Goodfellow et al., 2014), and diffusion models (Ho et al., 2020), have improved synthetic trace generation, these methods typically only generate specific fields, such as the number of requests or resource types (Bergsma et al., 2021), or are confined to fixed-structure traces, like network packets (Jiang et al., 2023; Yin et al., 2022).
@@
In this paper, we show how to use large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023) to generate synthetic workload traces. LLMs have been successfully adapted beyond natural language to domains like protein sequences (Shen et al., 2024), code (Roziere et al., 2023), and tabular data (Borisov et al., 2023). In addition, LLMs can produce outputs well-aligned with user inputs through fine-tuning (Ouyang et al., 2022; Wei et al., 2021) and generalize to new prompts at inference (Chung et al., 2022; Sanh et al., 2021). Thus, we posit that LLMs have the potential to generate synthetic traces that accurately capture real-world systems trace structures while adhering to user specifications.
@@
Despite their potential, using LLMs for synthetic systems trace generation presents significant challenges. Traces are often logged in tabular format and structured as graphs, which can vary in depth and width. Representing these traces as text sequences, optimal for modern autoregressive LLMs, is non-trivial. Moreover, trace data often contain complex implicit constraints that rely on relationships between multiple trace features. For example, an application process’s start time must precede that of all its child processes, while the parent process’s end time must be later than the end time of its children; such constraints need to hold across all nodes in the application’s graphical representation.
@@
Our approach adapts general-purpose LLMs to generate synthetic system traces, particularly focusing on microservice call graphs. Microservices are the de facto approach to designing modern cloud-based applications and power many popular services today (Uber, ; Netflix, ; Amazon, ). Microservice call graphs trace the execution of user requests through such applications. These call graphs form a type of trace with a rich directed acyclic graph (DAG) structure. We represent these graphs in a text-based format suitable for LLMs, enabling their use in trace generation. Crucially, one of our key innovations is to generate call graphs with structural constraints by recursively generating subgraphs, or layers. This approach allows the model to break down the complex task of reasoning about hierarchical graph structures and complex constraints into multiple simpler tasks. To further enhance the model’s ability to follow structural constraints and to meet user-requested attributes, we employ instruction tuning. During this phase, the model learns to explicitly generate a series of intermediate instructions between recursive layer generation steps, performing arithmetic and logical checks to ensure strict adherence to the desired structure.
@@
We validate our approach by fine-tuning Llama-2 7B (Touvron et al., 2023) on microservice trace data and conducting extensive evaluations. Results show that recursive generation with intermediate instructions significantly improves the model’s ability to generate valid outputs for complex call graphs. Compared to generative and probabilistic models, our synthetic traces better align with real trace distributions. Furthermore, we demonstrate that synthetic traces effectively replace real data for training microservice management tasks and that our fine-tuned model excels Llama-3.1 405B (Dubey et al., 2024) in trace feature prediction. We release our codes in https://github.com/ldos-project/TraceLLM.
@@ 2 Background
Figure 1: A simple social network application consists of eight microservices (Huye et al., 2023). Each user request triggers a sequence of microservice calls, forming a microservice call graph. The red lines represent the microservice call graph for a user request. Microservice call graphs are commonly logged in a tabular format, as shown in the figure on the right. Each row in the table represents a communication between two microservices with features in columns.
@@
Microservice Call Graphs.  In modern software architecture, an application is typically constructed as a constellation of multiple microservices (Gan et al., 2019; Luo et al., 2022; Huye et al., 2023), each with specific functionalities and dependencies on one another. When a user interacts with an application, such as sending an HTTP request, a complex sequence of communications among these microservices is triggered. Thus, a user request induces a microservice call graph, which maps the control/data flow and dependencies among the microservices involved in fulfilling the user’s request.
@@
Figure 1 illustrates a social network application with eight microservices. Red arrows indicate communications between microservices involved in processing the user’s request. The request first reaches a microservice (e.g., “Front end” in Figure 1) and waits for the communication to terminate. If the microservice requires additional communication to handle the request, then it triggers another microservice call (e.g., from “Front end” to “Authentication” in Figure 1). The communications triggered by a user’s request form a microservice call graph with four microservices. The vertices of the graph correspond to microservices (or the client), while the edges correspond to API calls invoking the microservices. Note that not all edges appear in the graph, as some services may not be invoked for a given request.
@@
A call graph can be represented as a tabular trace capturing API call features (i.e., edges) such as request source/destination, request type (e.g., HTTP and RPC), and start/finish times. Given their hierarchical structure, the tabular trace should preserve the parent-child relationships by ensuring that the child’s source matches the parent’s destination. Also, the start/end times of each call should be consistent with each other: (1) a microservice’s start time must precede its finish time, and (2) the parent-child relationships must be honored, i.e., the parent’s start (finish) time must precede (follow) the child’s. Finally, the IDs within a call graph (dot-decimal numbers provided for each call) must also be hierarchically connected to form a DAG.
@@
Synthetic Trace Generation using Machine Learning.  Microservice traces play a pivotal role in designing and evaluating techniques for improving the performance and reliability of microservice applications and optimizing the use of underlying resources. Representative use cases include techniques for critical path analysis (Zhang et al., 2022), anomaly detection (Xie et al., 2023), root cause analysis (Ikram et al., 2022), cluster management (Qiu et al., 2020), and cluster scheduling (Singhvi et al., 2021). Unfortunately, obtaining diverse real-world traces to study such techniques thoroughly remains challenging as publicly available traces are typically limited in size and only cover specific narrow settings far from the diversity expected when operating in the cloud.
@@
Given the importance and scarcity of public computer system traces, including microservice traces, recent studies have explored generative models for synthetic trace generation. Existing works (Lin et al., 2020; Jiang et al., 2023) leverage GAN (Goodfellow et al., 2014) and diffusion (Ho et al., 2020) models to generate network packet traces, while other work (Bergsma et al., 2021) uses LSTMs (Sherstinsky, 2020) to generate virtual machine workload traces. While these generative models are effective in their domains, the methods are limited to predicting specific fields or following training data distributions without conforming to structural constraints. These methods do not apply to microservice call graphs which requires handling hierarchical structures.
@@
Since traces are structured and can be represented in tabular form, machine learning methods for synthetic tabular data generation could be applied to synthetic trace generation. Recent approaches, such as TVAE (Xu et al., 2019) and GReaT (Borisov et al., 2023), leverage VAE (Kingma and Welling, 2013) and language models to advance synthetic tabular data generation techniques. However, these methods do not account for the hierarchical structure of call graphs within tabular representations. We provide a detailed comparison in § 4.
3 Training LLMs for Microservice Traces
@@
Our goal is to train a generative model for microservice call graph traces, enabling end-users to simulate diverse scenarios like rare microservice invocations that exhibit high response times. To achieve this, we condition the model’s output on user-specified attributes, including the invoked application, the number of microservice communications (i.e., graph edges), and overall application latency. Given the limitations of existing trace generation methods, we leverage LLMs. We initialize our model from LLMs pre-trained on large text datasets, as these models have proven effective when adapted to specialized domains such as proteins (Shen et al., 2024) and code (Roziere et al., 2023). Moreover, LLMs support flexible conditioning mechanisms, including natural language prompts (Ouyang et al., 2022) and structured input sequences (Borisov et al., 2023).
This section describes our approach to training an LLM for microservice call graph generation in two stages. First, we pre-train the model to capture the interactions in real-world call graph data and introduce a recursive subgraph generation to enhance its ability to generate large, structured graphs. Second, we fine-tune the model with instruction tuning, enabling it to generate call graphs with user-specified attributes and ensuring constraint adherence through natural language reasoning.
@@ 3.1.1 Encoding Call Graphs as Text
LLMs expect sequences of text as input, so we must encode our dataset of call graphs into text-based representations before training our model. As detailed in § 2 and shown in Figure 1, microservice call graphs are initially stored as tables. Rows represent edges (i.e., communications between microservices), while columns describe features for each edge. We follow the method proposed by GReaT (Borisov et al., 2023) and encode features in a natural language format. Our encoding procedure preserves all necessary information to recover the unique graph that produced the tabular data. Besides edge features, we also encode global attributes of the call graph to serve as conditioning information for the model.
A tabular call graph $\mathbf{X}$ has $m$ feature columns $\{f_{1},\dots,f_{m}\}$ and $n$ edge rows $\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$, where the value of feature $j$ for edge $i$ is $v_{ij}$. Each edge $\mathbf{x}_{i}$ is encoded as a text sequence $\mathbf{t}_{i}=[t_{i1},\dots,t_{im}]$, where $t_{ij}=[\phi(f_{j}),v_{ij}]$. Here, $\phi(f)$ converts feature name $f$ into a natural language template describing $v_{ij}$. For example, the encoding for edge 1 in Figure 1 would be: [Edge ID is 0, Source is Client, Destination is Front end, Type is HTTP, Communication starts at 0 ms, Communication finishes at 24 ms]. The full graph is represented as $\mathbf{t}=[\mathbf{t}_{1},\dots,\mathbf{t}_{n}]$, a sequence of text-encoded edges. Since call graph structure depends on feature values and not column order, we randomly shuffle feature order within each edge during training Borisov et al. (2023) to eliminate spurious position-based associations.
The overall call graph can be described by attributes such as maximum depth, total edges, and total communication latency. These attributes summarize complex interactions and serve as prompts for call graph generation. Let call graph $\mathbf{X}$ have $r$ attributes with names $\{a_{1},\dots,a_{r}\}$ and values $\{w_{1},\dots,w_{r}\}$. We encode them as a text sequence $\mathbf{c}=[c_{1},\dots,c_{r}]$, where $c_{j}=[a_{j},``:",w_{j}]$. See the Conditions in red in Figure 2 for an example. Attributes are prepended to each text-encoded call graph and predicted alongside edges during pre-training. Like edge features, attributes are randomly shuffled, and each is independently dropped with probability $p_{drop}$ to enable flexible prompting.